Question #4: Suppose Home is a small country. Use the graphs below to answer the questions.(a) Home Market Price $14 S 9 8 4 D 2 4 5 6 8 Quantity (b) Import Market Price X* +t $8 6 6 M 2 6 Import

a, Calculate Home consumer surplus and producer surplus in the absence of trade.

b, now suppose that Home engages in trade and faces the world price, P* = $6. Determine the consumer and producer surplus under free trade. Does Home benefit from trade? Explain.

c. Concerned about the welfare of the local producers, the Home government imposes a tariff in the amount of $2 (i.e., t = $2). Determine the net effect of the tariff on the Home economy.
Answer #1

Part a

Consumer Surplus

= 1/2 * ( 14 - 9) * (5 - 0)

= 1/2 * 5 * 5 = $12.5

Producer Surplus

= 1/2 * ( 9 - 4 ) * (5 - 0)

= 1/2 * 5 * 5 = $12.5

Total Surplus without trade = $12.5 + $12.5 = $25

Part b

Consumer Surplus

= 1/2 * ( 14 - 6) * (8 - 0)

= 1/2 * 8 * 8

= $32

Producer Surplus

= 1/2 * ( 6-4) * ( 2 - 0)

= $2

Total Surplus with trade

= $32 + $2 = $34

Home is better with trade as there is increase of $9.

Part c

Consumer Surplus

= 1/2 * ( 14 - 8) * ( 6 - 0)

= $18

Producer Surplus

= 1/2 * ( 8 - 4) * (4-0)

= $8

Revenue Received by Government

Area of Rectangle with measurements as

= (8 - 6) * ( 6 - 4)

= 2 * 2 = $4

Net Effect

Decrease in Consumer Surplus = $32 - $18 = $14

Rise in Producer Surplus = $8 - $2 = $6

Rise in Government Revenue = $4

Net Effect = 6 + 4 - 14 = -$4